OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -3, -12, -48, -192, ... a. an = 4 • -3n + 1 b. an = -3 • 4n - 1 c. an = -3 • 4n d. an = 4 • -3n

OpenStudy (anonymous):

does anyone know how to do this??

OpenStudy (anonymous):

this is an geometric progression with a1= -3 and r=-12/-3 =4 now formula for the nth term of a geometric series is \[a _{n}=a _{1} .r ^{n-1}\] now you can calculate......

OpenStudy (anonymous):

so option b would be right??

OpenStudy (anonymous):

and your options should be a. an = 4 • -3^(n + 1) b. an = -3 • 4^(n - 1) c. an = -3 • 4^n d. an = 4 • -3^n so b

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you are welcome @thinktallthoughts